A rookie quarterback throws a football with an initial upward velocity component of 12.0 m/s and a horizontal component of 20.0 m/s. How much time is required for the football to reach the highest point of its trajectory? How high is this point?
1
Expert's answer
2017-10-13T15:15:07-0400
v_y=gt
The time is
t=v_y/g=12.0/9.81=1.22 s.
The height is

h=v_y^2/2g=(12)^2/(2(9.81))=7.34 m.